OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 8x6 – 56x3 + 98. Provide the factors.

OpenStudy (anonymous):

2(2x^3+7)^2 So what we have here if we work from factored form out is: 2*(2x^3+7)(2x^3+7) Ignore the outside 2 for now. 2x^3 * 2x^3 = 4x^6 2x^3 * 7 = 14x^3 7 * 2x^3 = 14x^3 7 * 7 = 49 So we have: 2 * (4x^6 + 14x^3 + 14x^3 + 49) Distributing the two gives us: 8x^6 + 2*(28x^3) + 98 which is: 8x^6 + 56x^3 + 98. Which is your original problem. I like to plug these into my calculator and work them backwards. Hope this helps.

OpenStudy (anonymous):

Thank u